# On Prior Distributions for Orthogonal Function Sequences [^1] August 21, 2025

### v1

@@ Introduction
Modeling a sequence of orthogonal functions is required in a variety of statistical problems. Examples include principal component analysis (PCA) for functional data (Margaritella et al., 2021; Suarez and Ghosal, 2017), interactive fixed effects for panel data (Bai, 2009) and Karhunen-Loève expansion for stochastic processes (Yao et al., 2005). Such models entail a sequence of orthogonal functions to achieve interpretable and efficient representations. Although a Bayesian framework needs to specify a prior distribution over the sequence of functions, introducing priors that respect the orthogonality constraint is nontrivial. For instance, existing Bayesian FPCA typically place priors without explicitly enforcing orthogonality (e.g., Suarez and Ghosal, 2017). While such formulations offer computational convenience, they may compromise interpretability and the efficiency of representation, which are central motivations for modeling with orthogonal functions in the first place.
@@
There is a rich literature on incorporating constraints into Bayesian models, particularly in the context of shape-constrained function estimation. For example, monotonicity and convexity constraints have been widely studied in Bayesian nonparametric regression (Shively et al., 2009; Lin and Dunson, 2014; Lenk and Choi, 2017), along with other structural constraints such as unimodality and bounded variation (Wheeler et al., 2017; Yu et al., 2023). While these works focus on constraints imposed on single functions, enforcing mutual orthogonality among multiple functions poses additional challenges due to the rapidly increasing number of pairwise constraints. Several approaches have been proposed to encourage diversity among function components. Plumlee and Joseph (2018) proposed constraint Gaussian process to ensure orthogonality of a single function and pre-specified multiple functions. Constraint relaxation methods (Duan et al., 2020; Matuk et al., 2022) introduce soft penalties to approximate orthogonality, but these typically rely on a fixed tuning parameter, which may require careful calibration and cannot be adaptively learned from the data. As related approaches, repulsive priors (Petralia et al., 2012; Xie and Xu, 2020) can promote separation between functions but do not directly ensure orthogonality in the function space. Other alternatives include orthogonal random Fourier features (Yu et al., 2016), which produce random orthogonal basis functions but do not preserve orthogonality when used to construct multiple functional components. Similarly, models defined on the Stiefel manifold (e.g., Jauch et al., 2025) provide a principled way to impose orthogonal constraints through matrix-valued priors, though these can be computationally demanding and are often impractical when the basis functions themselves are not orthogonal. Post-hoc approaches that modify posterior samples to satisfy constraints (e.g., Sen et al., 2018) offer a practical alternative, but still lack theoretical guarantees.
@@
In this work, we propose the adaptive orthogonal prior (AOP), a novel prior distribution that directly induces approximate orthogonality among functions sequentially, in a computationally efficient and flexible manner. Our paper builds upon the framework of basis function expansions, where each function is represented as a linear combination of pre-specified basis functions. This formulation provides a flexible and interpretable structure that has been widely adopted in Bayesian modeling (e.g. Lang and Brezger, 2004) and offers a unified representation of functions, whose infinite-dimensional limit is known as Gaussian processes (Kimeldorf and Wahba, 1970; Williams and Rasmussen, 2006). Based on the basis function representation, AOP sequentially defines the prior distributions for the coefficient vectors of each function using conditionally Gaussian distributions with approximate orthogonality constraints. By controlling the strength of these constraints through hyperparameters, our prior flexibly accommodates both exact and approximate orthogonality, allowing researchers to balance strictness and model complexity depending on the application. AOP can be seamlessly incorporated into Bayesian functional principal component analysis (FPCA), where orthogonality among principal component functions is crucial for both interpretability and dimensionality reduction. Our prior not only guarantees nearly orthogonal posterior estimates of the principal functions but also enables automatic selection of the effective number of components owing to the sequential construction of AOP, leading to parsimonious low-rank representations. Furthermore, the conditional Gaussian formulation allows for straightforward implementation using standard Gibbs sampling, making posterior inference computationally tractable even in higher-dimensional settings.
@@
This paper is organized as follows. In Section 2, we introduce the AOP and discuss its theoretical properties. Section 3 presents the Bayesian FPCA model incorporating the proposed prior. We demonstrate the numerical performance of the proposed method through simulation studies in Section 4 and an application to human mobility data in Tokyo in Section 5. Finally, concluding remarks are given in Section 6. The R code implementing the proposed method is available at GitHub repository (https://github.com/sshonosuke/AOP).
@@
We then sequentially define the joint prior of $(\beta_{1},\ldots,\beta_{K})$. First, the prior of $\beta_{1}$ is defined as $\beta_{1}\sim{\mathcal{N}}_{L}(b_{01},B_{01})$ without any restrictions, where $b_{01}$ and $B_{01}$ are hyperparameters and ${\mathcal{N}}_{L}(\mu,\Sigma)$ denotes the $L$-dimensional normal distribution with mean vector $\mu$ and variance-covariance matrix $\Sigma$. Next, the prior of $\beta_{2}$ is defined, conditionally on $\beta_{1}$, as $\beta_{1}^{\top}\Omega\beta_{2}\sim{\mathcal{N}}(0,\tau_{2}^{2})$ and $H_{2}\beta_{2}\sim{\mathcal{N}}_{L}(b_{02},B_{02})$, assuming prior independence between $\beta_{1}^{\top}\Omega\beta_{2}$ and $H_{2}\beta_{2}$, where $b_{02}$, $B_{02}$ and $\tau_{2}^{2}$ are hyperparameters and $H_{2}$ is $(L-1)\times L$ matrix with rank L1L-1 that is independent of $\beta_{1}$. Note that the constraint (1) exactly holds when $\tau_{2}=0$. This conditional prior specification is equivalent to assigning a multivariate normal prior for $A_{2}(\beta_{1})\beta_{2}$ given $\beta_{1}$, where $A_{2}(\beta_{1})=((\beta_{1}^{\top}\Omega)^{\top},H_{2}^{\top})^{\top}$, that is, the first row vector of $A_{2}(\beta_{1})$ is $\beta_{1}^{\top}\Omega$ and the other part is equivalent to $H_{2}$. Hence, the conditional prior of $\beta_{2}$ given $\beta_{1}$ is given by
@@
While the proposed prior can be applied with fixed values of $\tau_{k}^{2}$, the specific choice of these hyperparameters can significantly influence the resulting posterior inference. In practice, strict orthogonality (i.e., $\tau_{k}^{2}=0$) may be desirable for leading functions to ensure interpretability, whereas milder constraints (i.e., larger $\tau_{k}^{2}$) for higher-order functions allow more flexibility and help capture finer structures in the data.
To accommodate data-driven adaptation of these constraints, we introduce a hierarchical approach by assigning prior distributions to $\tau_{k}^{2}$. Specifically, we consider two strategies: a global prior and a local prior. In the global prior approach, we assume a common strength of constraint across all components, that is, $\tau_{k}^{2}=\tau^{2}$ for all $k$, and assign an inverse-gamma prior $\tau^{2}\sim\mathrm{IG}(a_{0},b_{0})$ with fixed hyperparameters $(a_{0},b_{0})$. This specification enforces a uniform level of orthogonality and is suitable when the same degree of orthogonality is desired across all functions. In contrast, the local prior approach allows component-specific adaptation by assuming $\tau_{k}^{2}\sim\mathrm{IG}(a_{0},b_{0})$ independently for each $k$. This formulation enables greater flexibility, allowing stronger constraints on leading functions while relaxing constraints on higher-order functions, which can be advantageous when capturing complex variations in the data. Because of such adaptivity of strength of orthogonal constraint, we call the prior adaptively orthogonal prior (AOP).
From a computational perspective, introducing priors for $\tau_{k}^{2}$ only adds an additional inverse-gamma update step in the Gibbs sampler, preserving the overall simplicity of the posterior computation. Regarding the hyperparameters $(a_{0},b_{0})$, we use the default setting $a_{0}=3$ and $b_{0}=2/K^{2}$, which implies that the prior mean of $\tau^{2}$ is $E[\tau^{2}]=1/K^{2}$. This choice provides a weakly informative prior that favors moderate orthogonality (as $\tau^{2}$ is shrunk around a small value), while still allowing sufficient flexibility to adapt to the data.
Properties of the AOP prior
In this section, we investigate properties of the proposed AOP prior and clarify the effect of the hyperparameters. For simplicity, we assume that $\Omega=I_{L}$, that is, the basis functions are orthonormal. Furthermore, we consider the choice of hyperparameter as $b_{0,j}=0_{L-j}$ and $B_{0,j}=\gamma I_{L-j}$ for some $\gamma>0$, and assume that the $(L-j)\times L$ matrix $H_{j+1}$ is orthogonal, that is, $H_{j+1}H_{j+1}^{\top}=I_{L-j}$. Under the setting, we consider the conditional expectation $E[\|\beta_{j+1}\|^{2}|\beta_{1:j}]$, the conditional expectation of the squared norm of the coefficient vector of $f_{j+1}(x)$. It follows that
@@
${\rm tr}$ $\{{\rm Var}(\beta_{j+1}|\beta_{1:j})\}$
$={\rm tr}\left\{{\rm blockdiag}(\tau_{j+1}^{2}I_{j},\gamma I_{L-j})(A_{j+1}A_{j+1}^{\top})^{-1}\right\}$
$=\tau_{j+1}^{2}{\rm tr}\{(B_{1:j}^{\top}P_{j}B_{1:j})^{-1}\}+\gamma(L-j)+\gamma\ {\rm tr}\left\{(B_{1:j}^{\top}P_{j}B_{1:j})^{-1}B_{1:j}^{\top}H_{L-j}^{\top}(I_{L}-P_{j})H_{L-j}\right\}.$
Since $E[\beta_{j+1}|\beta_{1:J}]=0_{L}$ from (2), the above expression is equivalent to $E[\|\beta_{j+1}\|^{2}|\beta_{1:j}]$. The conditional expectation of the squared norm (conditional variance) is an increasing function of $\tau_{j}^{2}$, indicating that weaker orthogonal conditions lead to larger conditional variance. Also, under fixed $\tau_{j}^{2}$ and $\gamma$, the conditional variance decreases as $j$ increases, showing that the coefficient $\beta_{j}$ for the $j$th function will be more concentrated around the origin for larger $j$. Such property would be preferable since the $j$th function with large $j$ would not have much contribution.
To visualize the proposed prior, we set $L=K=2$ and $\beta_{1}=(0.5,1)$, and consider the conditonal prior for $\beta_{2}$. In Figure 1, we show the contours of the conditional densities under three choices of the hyperparameters. We observe that when the orthogonality constraint is strong (e.g., $\tau_{2}^{2}=0.01$), the conditional prior distribution is tightly concentrated along the line orthogonal to $\beta_{1}^{\top}\Omega\beta_{2}=0$, effectively enforcing near-orthogonality between the two functions. In contrast, a weaker constraint (e.g., $\tau_{2}^{2}=1$) leads to a broader prior distribution, allowing for more flexibility but reducing orthogonality. Additionally, increasing the variance $B_{02}$ results in a more diffuse prior marginally, while still respecting the constraint level imposed by $\tau_{2}^{2}$. These visualizations highlight how the proposed prior adaptively balances orthogonality and variability through its hyperparameters.
Figure 1: Conditional prior of $\beta_{2}$ given $\beta_{1}$ with $(\tau_{2}^{2},B_{02})=(0.01,1)$ (black), $(1,1)$ (red) and $(0.01,2)$ (blue). The solid line corresponds to the parameter space where the orthogonal condition holds exactly.
@@
We provide a posterior computation algorithm for the proposed prior. We first note that the joint prior density of $\beta_{1},\ldots,\beta_{K}$ can be expressed as
$\prod_{j=2}^{K}\exp\left\{-\frac{1}{2\tau_{j}^{2}}\sum_{k=1}^{j-1}(\beta_{k}^{\top}\Omega\beta_{j})^{2}\right\}\times\prod_{j=1}^{K}\exp\left\{-\frac{1}{2}(H_{j}\beta_{j}-b_{0j})^{\top}B_{0j}^{-1}(H_{j}\beta_{j}-b_{0j})\right\},$
@@
$\exp\left\{-\frac{1}{2\tau_{k}^{2}}\sum_{j=1}^{k-1}(\beta_{j}^{\top}\Omega\beta_{k})^{2}-\frac{1}{2}\sum_{j=k+1}^{K}\frac{(\beta_{j}^{\top}\Omega\beta_{k})^{2}}{\tau_{j}^{2}}-\frac{1}{2}(H_{k}\beta_{k}-b_{0k})^{\top}B_{0k}^{-1}(H_{k}\beta_{k}-b_{0k})\right\},$
which is a multivariate normal distribution ${\mathcal{N}}_{L}(A_{\beta_{k}}^{-1}B_{\beta_{k}},A_{\beta_{k}}^{-1})$, where
$A_{\beta_{k}}=H_{k}^{\top}B_{0k}^{-1}H_{k}+\frac{1}{\tau_{k}^{2}}\sum_{j=1}^{k-1}\Omega\beta_{j}\beta_{j}^{\top}\Omega+\sum_{j=k+1}^{K}\frac{\Omega\beta_{j}\beta_{j}^{\top}\Omega}{\tau_{j}^{2}},$
$B_{\beta_{k}}=H_{k}^{\top}B_{0k}^{-1}b_{0k}.$
Hence, when the likelihood function for $\beta_{k}$ is Gaussian, we can use a simple Gibbs sampling to generate random draws from the posterior. The specific example will be given in Section 3.
@@
where $f_{k}(t)$ is a principal function common to all the observations, $Z_{ik}$ is an observation-specific random weight for that function and ${\varepsilon}_{i}(t)$ corresponds to an error term. We assume that $Z_{ik}\sim{\mathcal{N}}(0,\lambda_{k})$, where $\lambda_{k}$ is an unknown parameter satisfying $\lambda_{1}>\cdots>\lambda_{K}$. Let $t_{i1},\ldots,t_{im_{i}}$ be a set of observed points for the $i$th observation. By introducing basis function representation for the principal function $f_{k}(x)$, Bayesian inference on FPCA is considered in Suarez and Ghosal (2017) and Nolan et al. (2025), and its robust version is also discussed in Zhang et al. (2025). However, the prior distributions for $f_{k}(x)$ does not guarantee orthogonality each other, which may lead to inefficient low-rank representations. We here introduce the proposed orthogonal prior for a sequence of functions, $f_{1}(x),\ldots,f_{K}(x)$, to obtain (nearly) orthogonal posterior inference on the principal functions.
@@
    (Sampling of $\beta_{k}$)   For $k=1,\ldots,K$, draw $\beta_{k}$ from its full conditional distribution, ${\mathcal{N}}_{L}(V_{k}^{-1}U_{k},V_{k}^{-1})$, where
@@
    Here, $\Phi_{i}=(\phi(t_{i1}),\ldots,\phi(t_{im_{i}}))^{\top}$ is the $m_{i}$-dimensional vector of basis functions evaluated at the observed points, and $F_{il}=(f_{l}(t_{i1}),\ldots,f_{l}(t_{im_{i}}))^{\top}$ denotes the principal function evaluated at the observed points.
@@
As shown above, all these full conditional distributions are standard distributions, which allows for efficient posterior computation without the need for complex sampling algorithm like rejection steps.
@@
$f_{1}(t)$ $=\sqrt{3}\,(2t-1),\ \ \ \ f_{2}(t)=\sqrt{5}\,(6t^{2}-6t+1),\ \ \ \ f_{3}(t)=\sqrt{7}\,(20t^{3}-30t^{2}+12t-1)$

### v2

# Sequential Adaptive Priors for Orthogonal Functions [^1] December 24, 2025

@@ Introduction
Modeling a sequence of orthogonal functions is necessary for a variety of statistical problems. Examples include functional principal component analysis (FPCA) (Margaritella et al., 2021; Suarez and Ghosal, 2017), interactive fixed effects for panel data (Bai, 2009) and Karhunen-Loève expansion for stochastic processes (Yao et al., 2005). Such models entail a sequence of orthogonal functions to achieve interpretable and efficient representations, especially for selecting an effective number of components from data. Although a Bayesian framework needs to specify a prior distribution over such a sequence of functions, introducing priors that respect orthogonality constraints is nontrivial. For instance, existing Bayesian FPCA typically place priors without explicitly enforcing orthogonality (e.g., Suarez and Ghosal, 2017). While such formulations offer computational convenience, they may compromise interpretability and efficiency of representation, which are central motivations for modeling with orthogonal functions in the first place.
@@
There is a rich literature on incorporating constraints into Bayesian models, particularly in the context of shape-constrained function estimation. For example, monotonicity and convexity constraints have been widely studied in Bayesian nonparametric regression (Shively et al., 2009; Lin and Dunson, 2014; Lenk and Choi, 2017), along with other structural constraints such as unimodality and bounded variation (Wheeler et al., 2017; Yu et al., 2023). While these works focus on constraints imposed on single functions, enforcing mutual orthogonality among multiple functions poses additional challenges due to the rapidly increasing number of pairwise constraints. Several approaches have been proposed to encourage diversity among function components. Plumlee and Joseph (2018) proposed constraint Gaussian process to ensure orthogonality of a single function and pre-specified multiple functions. Constraint relaxation methods (Duan et al., 2020; Matuk et al., 2022) introduce soft penalties to approximate orthogonality, but these typically rely on a fixed tuning parameter, which may require careful calibration and cannot be adaptively learned from data. As related approaches, repulsive priors (Petralia et al., 2012; Xie and Xu, 2020) can promote separation between functions but do not directly ensure orthogonality in the function space. Other alternatives include orthogonal random Fourier features (Yu et al., 2016), which produce random orthogonal basis functions but do not preserve orthogonality when used to construct multiple functional components. Similarly, models defined on the Stiefel manifold (e.g., Jauch et al., 2025) provide a principled way to impose orthogonal constraints through matrix-valued priors, though these can be computationally demanding and are often impractical when the basis functions themselves are not orthogonal. Post-hoc approaches that modify posterior samples to satisfy constraints (e.g., Sen et al., 2018) offer a practical alternative, but still lack theoretical guarantees.
@@
In this work, we propose the Adaptive Orthogonal (AO) prior, a novel prior distribution that directly induces approximate orthogonality among functions sequentially, in a computationally efficient and flexible manner. Our paper builds upon the framework of basis function expansions, where each function is represented as a linear combination of pre-specified basis functions. This formulation provides a flexible and interpretable structure that has been widely adopted in Bayesian modeling (e.g. Lang and Brezger, 2004) and offers a unified representation of functions, whose infinite-dimensional limit is known as Gaussian processes (Kimeldorf and Wahba, 1970; Williams and Rasmussen, 2006). Based on the basis function representation, our AO prior sequentially defines prior distribution for the coefficient vectors of each function using conditionally Gaussian distributions with approximate orthogonality constraints. By controlling the strength of these constraints through hyperparameters, our prior flexibly accommodates both exact and approximate orthogonality, allowing researchers to balance strictness and model complexity depending on the application. AO prior can be seamlessly incorporated into Bayesian functional principal component analysis (FPCA), where orthogonality among principal component functions is crucial for both interpretability and dimensionality reduction. Our prior not only guarantees nearly orthogonal posterior estimates of the principal functions but also enables automatic selection of the effective number of components owing to the sequential construction of AO prior, leading to parsimonious low-rank representations. Furthermore, the conditional Gaussian formulation allows for straightforward implementation using standard Gibbs sampling, making posterior inference computationally tractable even in higher-dimensional settings.
@@
This paper is organized as follows. In Section 2, we introduce the AO prior and discuss its theoretical properties. Section 3 presents Bayesian FPCA model incorporating the proposed prior. We demonstrate numerical performance of the proposed method through simulation studies in Section 4 and an application to human mobility data in Tokyo in Section 5. Finally, concluding remarks are given in Section 6. The R code implementing the proposed method is available at GitHub repository (https://github.com/sshonosuke/AOP).
@@
We then sequentially define the joint prior of $(\beta_{1},\ldots,\beta_{K})$. First, the prior of $\beta_{1}$ is defined as $\beta_{1}\sim{\mathcal{N}}_{L}(b_{01},B_{01})$ without any restrictions, where $b_{01}$ and $B_{01}$ are hyperparameters and ${\mathcal{N}}_{L}(\mu,\Sigma)$ denotes the $L$-dimensional normal distribution with mean vector $\mu$ and variance-covariance matrix $\Sigma$. Next, the prior of $\beta_{2}$ is defined, conditionally on $\beta_{1}$, as $\beta_{1}^{\top}\Omega\beta_{2}\sim{\mathcal{N}}(0,\tau_{2}^{2})$ and $H_{2}\beta_{2}\sim{\mathcal{N}}_{L-1}(b_{02},B_{02})$, assuming prior independence between $\beta_{1}^{\top}\Omega\beta_{2}$ and $H_{2}\beta_{2}$, where $b_{02}$, $B_{02}$ and $\tau_{2}^{2}$ are hyperparameters and $H_{2}$ is $(L-1)\times L$ matrix with rank L1L-1 that is independent of $\beta_{1}$. Note that the constraint (1) exactly holds when $\tau_{2}=0$. This conditional prior specification is equivalent to assigning a multivariate normal prior as $A_{2}(\beta_{1})\beta_{2}|\beta_{1}\sim{\mathcal{N}}(b_{02}^{\ast},{\rm blockdiag}(\tau_{2}^{2},B_{02}))$ where $b_{02}^{\ast}=(0,b_{02}^{\top})^{\top}$ and $A_{2}(\beta_{1})=((\beta_{1}^{\top}\Omega)^{\top},H_{2}^{\top})^{\top}$. Note that the first row vector of $A_{2}(\beta_{1})$ is $\beta_{1}^{\top}\Omega$ and the other part is equivalent to $H_{2}$. Then, the conditional prior of $\beta_{2}$ given $\beta_{1}$ is given by
@@
While the proposed prior can be applied with fixed values of $\tau_{k}^{2}$, specific choice of these hyperparameters can significantly influence the resulting posterior inference. In practice, strict orthogonality (i.e., $\tau_{k}^{2}=0$) may be desirable for leading functions to ensure interpretability, whereas milder constraints (i.e., larger $\tau_{k}^{2}$) for higher-order functions allow more flexibility and help capture finer structures in the data.
To accommodate data-driven adaptation of these constraints, we introduce a hierarchical approach by assigning prior distributions to $\tau_{k}^{2}$. Specifically, we consider two strategies: a global prior and a local prior. In the global prior approach, we assume a common strength of constraint across all components, that is, $\tau_{k}^{2}=\tau^{2}$ for all $k$, and assign an inverse-gamma prior $\tau^{2}\sim\mathrm{IG}(a_{0},b_{0})$ with fixed hyperparameters $(a_{0},b_{0})$. This specification enforces a uniform level of orthogonality and is suitable when the same degree of orthogonality is desired across all functions. In contrast, the local prior approach allows component-specific adaptation by assuming $\tau_{k}^{2}\sim\mathrm{IG}(a_{0},b_{0})$ independently for each $k$. This formulation enables greater flexibility, allowing stronger constraints on leading functions while relaxing constraints on higher-order functions, which can be advantageous when capturing complex variations in the data. Because of such adaptivity of strength of orthogonal constraints, we call the prior Adaptively Orthogonal (AO) prior.
From a computational perspective, introducing priors for $\tau_{k}^{2}$ only adds an additional inverse-gamma update step in the Gibbs sampler, preserving the overall simplicity of the posterior computation. Regarding the hyperparameters $(a_{0},b_{0})$, we use the default setting $a_{0}=3$ and $b_{0}=2/K^{2}$, which implies that the prior mean of $\tau^{2}$ is $\mathbb{E}[\tau^{2}]=1/K^{2}$. This choice provides a weakly informative prior that favors moderate orthogonality (as $\tau^{2}$ is shrunk around a small value), while still allowing sufficient flexibility to adapt to the data.
Properties of the AO prior
In this section, we investigate theoretical properties of the proposed AO prior and clarify the effect of its hyperparameters. For simplicity, we assume that $\Omega=I_{L}$, that is, the basis functions are orthonormal. Furthermore, we consider the choice of hyperparameter as $b_{0,j}=0_{L-j}$ and $B_{0,j}=\gamma I_{L-j}$ for some $\gamma>0$, and assume that the $(L-j)\times L$ matrix $H_{j+1}$ is orthogonal, that is, $H_{j+1}H_{j+1}^{\top}=I_{L-j}$. Under this setting, we consider a conditional expectation $\mathbb{E}[\|\beta_{j+1}\|^{2}|\beta_{1:j}]$, a conditional expectation of the squared norm of the coefficient vector of $f_{j+1}(x)$. It follows that
@@
${\rm tr}$ $[{\rm Var}(\beta_{j+1}|\beta_{1:j})]$
$={\rm tr}\left[{\rm blockdiag}(\tau_{j+1}^{2}I_{j},\gamma I_{L-j})(A_{j+1}A_{j+1}^{\top})^{-1}\right]$
$=\tau_{j+1}^{2}{\rm tr}[(B_{1:j}^{\top}P_{j}B_{1:j})^{-1}\}+\gamma(L-j)+\gamma\ {\rm tr}\left\{(B_{1:j}^{\top}P_{j}B_{1:j})^{-1}B_{1:j}^{\top}H_{L-j}^{\top}(I_{L}-P_{j})H_{L-j}\right].$
Since $\mathbb{E}[\beta_{j+1}|\beta_{1:J}]=0_{L}$ from (2), the above expression is equivalent to $\mathbb{E}[\|\beta_{j+1}\|^{2}|\beta_{1:j}]$. The conditional expectation of the squared norm (conditional variance) is an increasing function of $\tau_{j}^{2}$, indicating that weaker orthogonal conditions lead to larger conditional variance. Also, under fixed $\tau_{j}^{2}$ and $\gamma$, the conditional variance decreases as $j$ increases, showing that the coefficient $\beta_{j}$ for the $j$th function will be more concentrated around the origin for larger $j$. Such property would be preferable since the $j$th function with large $j$ would not have much contribution.
To visualize the proposed prior, we set L=K=2L\!=\!K\!=\!2 and consider the conditional prior for $\beta_{2}$ given $\beta_{1}$. In Figure 1, we show the contours of the conditional densities under $\beta_{1}=(0.5,1),\tau_{2}^{2}=0.05$ and $\beta_{1}=(0.5,-1),\tau_{2}^{2}=0.3$, and that of the standard normal prior, $\beta_{2}\sim{\mathcal{N}}(0,10I_{2})$ for comparison, where $b_{02}=(0,0)^{\top}$ and $B_{02}=2I_{2}$ for the proposed prior. We observe that the proposed conditional prior has significant prior mass around the parameter space where the orthogonal constraint is exactly satisfied, i.e. $\{\beta_{2}=(t,-t/2),\ \ t\in(-\infty,\infty)\}$ for $\beta_{1}=(0.5,1)$ and $\{\beta_{2}=(t,-t/2),\ t\in(-\infty,\infty)\}$ $\beta_{2}=\{(t,t/2),\ t\in(-\infty,\infty)\}$. Moreover, the degree of concentration is dictated by the scale parameter $\tau_{2}$, so the smaller value of $\tau_{2}$ leads to more concentration on the orthogonal parameter space. Such properties are quite different from the standard independent prior, where the information of $\beta_{1}$ is not reflected to the prior of $\beta_{2}$. These visualizations highlight how the proposed prior adaptively balances orthogonality and variability through its hyperparameters.
Figure 1: Contour plots of AO prior for $\beta_{2}$ given $\beta_{1}$ with $\beta_{1}=(0.5,1),\tau_{2}^{2}=0.05$ (left) and $\beta_{1}=(0.5,-1),\tau_{2}^{2}=0.3$ (center) and the standard prior, $\beta_{2}\sim N(0,10I_{2})$ (right). The solid blue line in right and center panels corresponds to the parameter space of $\beta_{2}$ where the orthogonal condition holds exactly.
@@
We provide a posterior computation algorithm for the proposed prior. From the conditional prior specification (2), we first note that the joint prior of $\beta_{1},\ldots,\beta_{K}$ can be obtained as
$\prod_{j=1}^{K-1}|{\rm det}(A_{j+1}(\beta_{1:j}))|(\tau_{j+1}^{2})^{-j/2}\exp\left[-\frac{1}{2\tau_{j+1}^{2}}\sum_{k=1}^{j}(\beta_{k}^{\top}\Omega\beta_{j+1})^{2}\right]$
$\times\prod_{j=1}^{K}\exp\left[-\frac{1}{2}(H_{j}\beta_{j}-b_{0j})^{\top}B_{0j}^{-1}(H_{j}\beta_{j}-b_{0j})\right],$
@@
$\prod_{j=k}^{K-1}|{\rm det}(A_{j+1}(\beta_{1:j}))|\exp\left[-\frac{1}{2\tau_{k}^{2}}\sum_{j=1}^{k-1}(\beta_{j}^{\top}\Omega\beta_{k})^{2}-\frac{1}{2}\sum_{j=k+1}^{K}\frac{(\beta_{j}^{\top}\Omega\beta_{k})^{2}}{\tau_{j}^{2}}\right]$
$\times\exp\left[-\frac{1}{2}(H_{k}\beta_{k}-b_{0k})^{\top}B_{0k}^{-1}(H_{k}\beta_{k}-b_{0k})\right].$
Then, without the term $\prod_{j=k}^{K-1}|A_{j+1}(\beta_{1:j})|$, the above full conditional is a multivariate normal distribution ${\mathcal{N}}_{L}(A_{\beta_{k}}^{-1}B_{\beta_{k}},A_{\beta_{k}}^{-1})$, where
$A_{\beta_{k}}=H_{k}^{\top}B_{0k}^{-1}H_{k}+\frac{1}{\tau_{k}^{2}}\sum_{j=1}^{k-1}\Omega\beta_{j}\beta_{j}^{\top}\Omega+\sum_{j=k+1}^{K}\frac{\Omega\beta_{j}\beta_{j}^{\top}\Omega}{\tau_{j}^{2}},\ \ \ \ \ B_{\beta_{k}}=H_{k}^{\top}B_{0k}^{-1}b_{0k}\,.$
Hence, when the likelihood function for $\beta_{k}$ is Gaussian, we can use a simple Metropolis-Hastings algorithm via generating proposals from the normal distribution. The specific example will be given in Section 3.
Sparsity-inducing prior
Since the AO prior is a valid multivariate prior distribution, we can introduce sparsity-inducing prior into the AO prior. Specifically, we adopt a hierarchical prior of the form, $H_{k}\beta_{k}\sim\mathcal{N}_{L-k+1}(0,\gamma_{k}I_{L-k+1})$ and $\sqrt{\gamma_{k}}\sim{\rm HC}_{+}$, where $\rm HC$$_{+}$ denotes the standard half-Cauchy prior. This corresponds to setting $B_{0k}=\gamma_{k}I_{L-k+1}$ in the general construction of the AO prior and this formulation leads to the horseshoe prior (Carvalho et al., 2009) for the marginal prior of $H_{k}\beta_{k}$. When $H_{k}\beta_{k}$ is close to the origin, the $k$th function is $f_{k}(x)\approx 0$, so that irrelevant functions could be automatically eliminated in the posterior. To perform the posterior computation with the sparsity-inducing prior, we can use the sampling algorithm in the previous section, under given $\gamma_{k}$. To generate random sample of $\gamma_{k}$, we introduce an additional variable $\eta_{k}$ to represent $\gamma_{k}|\eta_{k}\sim{\rm IG}(1/2,1/\eta_{k})$ and $\eta_{k}\sim{\rm IG}(1/2,1)$, which gives $\sqrt{\gamma_{k}}\sim{\rm HC}_{+}$ as the marginal distribution. Then, sampling from the full conditional of $\gamma_{k}$ is performed by two steps. We first generate $\eta_{k}$ from ${\rm IG}(1,1+1/\gamma_{k})$ and then generate $\gamma_{k}$ from ${\rm IG}(\widetilde{a}_{k},\widetilde{b}_{k})$, where
$\widetilde{a}_{k}=\frac{1}{2}+\frac{L-k}{2},\ \ \ \ \ \widetilde{b}_{k}=\frac{1}{2}(H_{k}\beta_{k}-b_{0k})^{\top}(H_{k}\beta_{k}-b_{0k})+\frac{1}{\eta_{k}}.$
@@
where $f_{k}(t)$ is a principal function common to all the observations, $Z_{ik}$ is an observation-specific random weight for that function and ${\varepsilon}_{i}(t)$ corresponds to an error term. We assume that $Z_{ik}\sim{\mathcal{N}}(0,\lambda_{k})$, where $\lambda_{k}$ is an unknown parameter satisfying $\lambda_{1}>\cdots>\lambda_{K}$. Let $t_{i1},\ldots,t_{im_{i}}$ be a set of observed points for the $i$th observation. By introducing basis function representation for the principal function $f_{k}(x)$, Bayesian inference on FPCA is considered in Suarez and Ghosal (2017) and Nolan et al. (2025), and its robust version is also discussed in Zhang et al. (2025). However, their prior distributions for $f_{k}(x)$ does not guarantee orthogonality each other, which may lead to inefficient low-rank representations. We here introduce the proposed orthogonal prior for a sequence of functions, $f_{1}(x),\ldots,f_{K}(x)$, to obtain (nearly) orthogonal posterior inference on the principal functions.
@@
    (Sampling of $\beta_{k}$) For $k=1,\ldots,K$, generate a proposal $\beta_{k}^{\ast}$ of $\beta_{k}$ from ${\mathcal{N}}_{L}(V_{k}^{-1}U_{k},V_{k}^{-1})$, where
@@
    Here, $\Phi_{i}=(\phi(t_{i1}),\ldots,\phi(t_{im_{i}}))^{\top}$ is the $m_{i}$-dimensional vector of basis functions evaluated at the observed points, and $F_{il}=(f_{l}(t_{i1}),\ldots,f_{l}(t_{im_{i}}))^{\top}$ denotes the principal function evaluated at the observed points. Then, accept the proposal $\beta_{k}^{\ast}$ with probability
    $\min\bigg(1,\prod_{j=k}^{K-1}\frac{|{\rm det}(A_{j+1}(\beta_{1:j}^{\ast}))|}{|{\rm det}(A_{j+1}(\beta_{1:j}))|}\bigg),$
    where $\beta_{1:j}^{\ast}$ is defined by replacing $\beta_{k}$ in $\beta_{1:j}$ with $\beta_{k}^{\ast}$.
@@
As shown above, all these full conditional distributions are standard distributions, which allows for efficient posterior computation.
@@
$f_{1}(t)$ $=\sqrt{3}\,\bigl(2t-1\bigr),~f_{2}(t)=\sqrt{5}\,\bigl(6t^{2}-6t+1\bigr),~f_{3}(t)=\sqrt{7}\,\bigl(20t^{3}-30t^{2}+12t-1\bigr)$